OpenStudy (anonymous):

Given the exponential equation 2x = 8, what is the logarithmic form of the equation in base 10?

OpenStudy (anonymous):

WAIT actually I might have got it, so could someone just check my answer?

OpenStudy (anonymous):

Is it log(10) 8 / log(10) 2????

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2^x=8}\) The relation is \(\normalsize\color{black}{ A^B=C~~~~~~~~→~~~~~~\log_AC=B}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2^x=8~~~~~~~→~~~~~?}\) you tell me

OpenStudy (anonymous):

log(2)8 = x

OpenStudy (solomonzelman):

Yes, but you need in base of 10, so apply, \(\normalsize\color{blue}{ \log_DC=\huge\frac{Log_{10}C}{Log_{10}D}}\)

OpenStudy (anonymous):

exactly, couldn't you apply base of 10 from the beginning?

OpenStudy (solomonzelman):

well, not really, because of the relation, you at first need to have the base 2.

OpenStudy (anonymous):

okay, so from there how do you go about switching it into base of 10?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I posed the rule, log(D) C = Log (10) C / Log (10) D

OpenStudy (solomonzelman):

*posted

OpenStudy (solomonzelman):

well, base 10 is just when the base is unspecified anyway

OpenStudy (anonymous):

how do you get D though?

OpenStudy (anonymous):

@SolomanZelman